Table 4-8 Firm Xs Firm Zs Firm Ys Quantity Quantity Quantity Supplied Price Supplied Supplied $0 $3 12 $6 14 12 18 $9 24 $
Answer #1

ANSWER:

When price is $12 , total quantity supplied by firm x and y is 8 + 16 = 24

When price is $15 , total quantity supplied by firm x and y is 10 + 20 = 30

Increase in quantity supplied = 30 - 24 = 6

so option c is the right answer.